OpenStudy (anonymous):

A line segment, ST, has endpoints S(-10,1) and T(-2,3). Which of the following equations represents the perpendicular bisector of the line segment? y = -4x - 22 y = 8x - 22 y = 2x - 23 y = -4x - 30

OpenStudy (anonymous):

im confused with this one

OpenStudy (anonymous):

first get the slope,wha do you think will be the slope?

OpenStudy (anonymous):

will it b -10,1

OpenStudy (anonymous):

i mean if you connect S and T,what is their slope?

OpenStudy (anonymous):

are you familiar on how to get slope between two points?

OpenStudy (anonymous):

S(-10,1) T(-2,3)

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

oh,i see are you familiar with m=(y2-y1)/(x2-x1) ?

OpenStudy (anonymous):

looks familiar but isnt that what the variables represent

OpenStudy (anonymous):

S and T are points so your y2 will be 3,your y1will be 1, your x2 will be -2,and your x1 will be -10

OpenStudy (anonymous):

try to substitute now

OpenStudy (anonymous):

so am i suppose to add or subtract

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

ok so this is for the problem im doing now

OpenStudy (anonymous):

because if i subtract all of that i will get 3

OpenStudy (anonymous):

what? that's not it,can you show how did you substitute?

OpenStudy (anonymous):

im sorry im just so confused im doing the problem and wht u sent me

OpenStudy (anonymous):

oky,do you like me to show you how?

OpenStudy (anonymous):

this problem right A line segment, ST, has endpoints S(-10,1) and T(-2,3). Which of the following equations represents the perpendicular bisector of the line segment? y = -4x - 22 y = 8x - 22 y = 2x - 23 y = -4x - 30

OpenStudy (anonymous):

that's the problem you want to solve

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i'll show the solution for getting the slope?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your slope for the line segment ST will be m=(3-1) / (-2-(-10)) m=2/(-2+10) m=2/8 m=1/4

OpenStudy (anonymous):

and i remove the one right and it leaves me with 4

OpenStudy (anonymous):

no,you would still negate it, so the slope of the other equation will be -4

OpenStudy (anonymous):

so its a negative instead do i do the second part the same way as the first

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

for perpendicular lines their slope would be negative reciprocal,so get its reciprocal and have a negative sign

OpenStudy (anonymous):

so it would be D -4 -30

OpenStudy (anonymous):

what do you think the slope of letter A?

OpenStudy (anonymous):

how are you sure with letter D?

OpenStudy (anonymous):

wait i got it thank u

OpenStudy (anonymous):

what did you got?